If Josie were born 4 years earlier, she would be 2 less than twice her age if she were born 4 years later. how old is she now?
Answer by fractalier(6550)

Call Josie's age NOW, x. Then we have
x + 4 = 2(x - 4) - 2
x + 4 = 2x - 8 - 2
x + 4 = 2x - 10
x = 2x - 14
x = 14 years old now